# arXiv:2602.23189v2[math.AP] 02 Mar 2026

### v2

@@ 2.2 First part of the entropy
We begin by estimating the time derivative of ${\mathcal{E}}_{1}$ for sufficiently smooth functions $(\rho_{+},\rho_{-},\alpha_{+},u)$ without assuming that they solve System (LABEL:system_eps).
@@
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition (2.1), then for all $(\rho_{+},\rho_{-},\alpha_{+},u)$ functions verifying the regularity properties of the limit solution, see Theorem 1.3, such that $p_{i}(\rho_{i})=\rho_{i}^{\gamma_{i}}$, for $i=+,$-, and $p_{+}(\rho_{+})=p_{-}(\rho_{-})$, we have
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)\leqslant$ $-\mu\int_{0}^{t}\int_{\Omega}\left(\nabla(u^{\varepsilon}-u)\right)^{2}-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}$
$+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\int_{\Omega}\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-})-\left(\alpha_{+}^{\varepsilon}p_{+}\left(\rho_{+}^{\varepsilon}\right)+\alpha_{-}^{\varepsilon}p_{-}\left(\rho_{-}^{\varepsilon}\right)\right)\Big)\operatorname{div}u$
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$
$-\mu\int_{0}^{t}\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u).$
@@
$\begin{split}{\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)=&E(t)-E(0)\\ &+\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(t,\cdot)-\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(0,\cdot) \\ &-\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)\\ &-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(t,\cdot)\\ &+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(0,\cdot).\end{split}$ (2.7)
@@
$\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)=-\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}u^{\varepsilon}\cdot(\partial_{t}u+(u^{\varepsilon}\cdot\nabla)u)\\ -\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u^{\varepsilon}:\nabla u+(\mu+\lambda)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}u^{\varepsilon}\right)(\operatorname{div}u),$
where we can replace
$-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$ $=-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon})\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$
$=-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+}^{\varepsilon})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}^{\varepsilon}))\operatorname{div}u.$
We can also rewrite
@@
$-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i}\Big)=\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}}\rho_{i}^{\gamma_{i}}=\int_{\Omega}\frac{1}{\varepsilon^{2}}\rho_{+}^{\gamma_{+}}$
hence
$\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i}\Big)(t,\cdot)+\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i})\Big)(0,\cdot)\\ =\int_{0}^{t}\!\!\!\int_{\Omega}\frac{1}{\varepsilon^{2}}(\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon})\partial_{t}\rho_{+}^{\gamma_{+}}=\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\frac{1}{\varepsilon^{2}}\alpha_{i}^{\varepsilon}\gamma_{i}\rho_{i}^{\gamma-1}\partial_{t}\rho_{i}$
$-\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\rho_{i}^{\gamma_{i}-1}\rho_{i}^{\varepsilon}(t,\cdot)+\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\rho_{i}^{\gamma_{i}-1}\rho_{i}^{\varepsilon}(0,\cdot)$ $=-\int_{0}^{t}\!\!\!\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}}{\varepsilon^{2}}\Big(\partial_{t}\rho_{i}+(u^{\varepsilon}\cdot\nabla)\rho_{i}\Big).$
We then have obtained
$-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(t,\cdot)$
$+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(0,\cdot)$ $=\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$
$-\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}\alpha_{i}\rho_{i}u\cdot\nabla\rho_{i}$
$-\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}\alpha_{i}\rho_{i}u\cdot\nabla\rho_{i}=$ $-\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\alpha_{i}u\cdot\nabla p_{i}(\rho_{i})=-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}u\cdot\nabla p_{+}(\rho_{+})$
$=$ $\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}p_{+}(\rho_{+})\operatorname{div}u$
$=$ $\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}))\operatorname{div}u$
where we have used that $u$ satisfies the Dirichlet boundary condition.
Therefore, using the energy estimate (2.4) satisfied by $(\rho_{+}^{\varepsilon},\rho_{-}^{\varepsilon},\alpha_{+}^{\varepsilon},u^{\varepsilon})$ and plugging all the previous computations into (2.7), we obtain:
[…]
@@
$-\operatorname{div}u^{\varepsilon}\operatorname{div}(u^{\varepsilon}-u)$ $=-\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}-\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u).$
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right):=\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$ (2.8a)
$+\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})$ (2.8b)
$+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\int_{\Omega}\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-})-\left(\alpha_{+}^{\varepsilon}p_{+}\left(\rho_{+}^{\varepsilon}\right)+\alpha_{-}^{\varepsilon}p_{-}\left(\rho_{-}^{\varepsilon}\right)\right)\Big)\operatorname{div}u$ (2.8c)
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$ (2.8d)
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$ (2.8e)
$-\mu\int_{0}^{t}\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u).$ (2.8f)
@@
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right)=$ $\mu\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}-\rho}{\rho}\Delta u(u-u^{\varepsilon})$
$+\int_{0}^{t}\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u)$
$-\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon}).$
@@
First, let us note that because we consider a solution of System (1.5), we have $\operatorname{div}u=0$, $\partial_{t}\rho_{i}=0$ and $\nabla\rho_{i}=0$, therefore (2.8c), (2.8d) and (2.8e) in the remainder $\mathcal{R}$} are cancelled and we only have
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right)=$ $\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$
$+\int_{0}^{t}\mu\int_{\Omega}\nabla u:\nabla(u-u^{\varepsilon}).$
Denoting $\rho=\alpha_{+}\rho_{+}+\alpha_{-}\rho_{-}$ and $\rho_{\varepsilon}=\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}$, we use the momentum equation verified by $(\rho_{+},\rho_{-},\alpha_{+},u)$ to get:
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right)=\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\mu\Delta u(u-u^{\varepsilon})+\mu\int_{0}^{t}\int_{\Omega}\nabla u:\nabla(u-u^{\varepsilon})\\ -\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi\cdot(u-u^{\varepsilon})+\int_{0}^{t}\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u).$
@@
2.3 Second part of the entropy
We now study the evolution of the entropy component (2.3) for weak solutions of System (LABEL:system_eps).

### v3

@@ 3.1 First part of the entropy
We begin by estimating the time derivative of ${\mathcal{E}}^{\varepsilon}_{1}$ for sufficiently smooth functions $(\rho_{+},\rho_{-},\alpha_{+},u)$ without assuming that they solve System (LABEL:system_eps).
@@
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition 2.1, then for all $(\rho_{+},\rho_{-},\alpha_{+},u)$ functions verifying the regularity properties of the limit solution, see Theorem 2.3, such that $p_{+}(\rho_{+})=p_{-}(\rho_{-})=C_{0}>0$ and $\operatorname{div}u=0$, we have
${\mathcal{E}}_{1}^{\varepsilon}(t)-{\mathcal{E}}_{1}^{\varepsilon}(0)\leqslant$ $-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\left(\nabla(u^{\varepsilon}-u)\right)^{2}-(\lambda+\mu)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}\,\mathrm{d}x\,\mathrm{d}s$
$-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)\,\mathrm{d}x\,\mathrm{d}s$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})\,\mathrm{d}x\,\mathrm{d}s$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})\,\mathrm{d}x\,\mathrm{d}s.$
@@
${\mathcal{E}}_{1}^{\varepsilon}(t)-{\mathcal{E}}_{1}^{\varepsilon}(0)=$ $E^{\varepsilon}(t)-E^{\varepsilon}(0)$
$+\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(t,\cdot)-\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(0,\cdot)$
$-\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot).$
@@
$\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)=-\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}u^{\varepsilon}\cdot(\partial_{t}u+(u^{\varepsilon}\cdot\nabla)u)\\ -\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u^{\varepsilon}:\nabla u+(\mu+\lambda)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}u^{\varepsilon}\right)(\operatorname{div}u).$
We can rewrite
@@
$+\int_{\Omega}\rho^{\varepsilon}\Big(((u^{\varepsilon}-u)\cdot\nabla)u\Big)\cdot(u-u^{\varepsilon}).$
Therefore, using the energy estimate (2.2) satisfied by $(\rho_{+}^{\varepsilon},\rho_{-}^{\varepsilon},\alpha_{+}^{\varepsilon},u^{\varepsilon})$, we obtain:
${\mathcal{E}}_{1}^{\varepsilon}(t)-{\mathcal{E}}_{1}^{\varepsilon}(0)\leqslant$ $-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\left(\nabla u^{\varepsilon}\right)^{2}\,\mathrm{d}x\,\mathrm{d}t-\left(\lambda+\mu\right)\int_{0}^{t}\!\!\!\int_{\Omega}(\operatorname{div}u^{\varepsilon})^{2}\,\mathrm{d}x\,\mathrm{d}t$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})$
$-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$
$+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u^{\varepsilon}:\nabla u+(\mu+\lambda)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}u^{\varepsilon}\right)(\operatorname{div}u).$
By virtue of monotonicity of the operators $\nabla$ and $\operatorname{div}$, namely, we write
@@
$-\operatorname{div}u^{\varepsilon}\operatorname{div}(u^{\varepsilon}-u)$ $=-\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}-\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u),$
which ends the proof because $u$ is divergence free. ∎
Remark 3.2.
Appendix A presents entropy computations for more general velocity fields $u$, which extend those considered here and could be used to study stability or weak-strong uniqueness.
@@
${\mathcal{E}}_{1}^{\varepsilon}(t)-{\mathcal{E}}_{1}^{\varepsilon}(0)\leqslant$ $-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\left(\nabla(u^{\varepsilon}-u)\right)^{2}-(\lambda+\mu)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}\,\mathrm{d}x\,\mathrm{d}s$
$+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\frac{\rho^{\varepsilon}-\rho}{\rho}\Delta u\cdot(u-u^{\varepsilon})+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u)$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\frac{\rho-\rho^{\varepsilon}}{\rho}\nabla\Pi\cdot(u-u^{\varepsilon})-\int_{0}^{t}\!\!\!\int_{\Omega}\nabla\Pi\cdot(u-u^{\varepsilon}).$
@@
${\mathcal{R}}^{\varepsilon}:=$ $-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon}).$
We use the momentum equation verified by $(\rho_{+},\rho_{-},\alpha_{+},u)$ to get:
${\mathcal{R}}^{\varepsilon}=\int_{0}^{t}\!\!\!\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\mu\Delta u(u-u^{\varepsilon})+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u:\nabla(u-u^{\varepsilon})\\ -\int_{0}^{t}\!\!\!\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi\cdot(u-u^{\varepsilon})+\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u).$
@@
3.2 Logarithmic entropies
We now study the evolution of the entropy component (1.8) for weak solutions of System (LABEL:system_eps). We begin by estimating the time derivative of ${\mathcal{E}}^{\varepsilon}_{2}$ for sufficiently smooth functions $(\rho_{+},\rho_{-},\alpha_{+},u)$ without assuming that they solve System (1.4).